# Why Human Guidance Matters in Collaborative Vibe Coding

### v1

@@ Results
Building on the previous results showing a strong advantage of humans as selectors and instructors, we next asked whether this advantage could be partially retained while reducing human involvement. Human labor is costly and difficult to scale, whereas AI can provide fast and generally comparatively less expensive automatic guidance. We next tested a hybrid human–AI vibe coding setting that mixes human and AI vibe coders within the same multi-round coding chain with the same and comparable code generation system.
Testing hybrid human-AI instructors and selectors
@@
In this experiment, we constructed hybrid coding chains in which selector and instructor roles were randomly assigned to either a human or an AI agent according to a fixed ratio (Fig. 5A). As in all other experiments, human and AI instructors each interacted with an AI code generation system to create the SVG code. We tested three hybrid conditions: human-75% (AI-25%, where humans take 75% of iterations and AI takes the rest), human-50% (AI-50%), and human-25% (AI-75%). For each condition, we repeated the 10 target images twice, resulting in 20 coding chains and 300 total iterations. Since we found that the results were already stable with one iteration, each human or AI-bot participated in only one iteration. We recruited 225 human participants for the human-75% condition, 150 for the human-50% condition, and 75 for the human-25% condition.
@@
As shown in Fig. 5B, all hybrid conditions performed better than fully AI-led vibe coding ($\Delta$=0.734, pp=0.007, Cohen’s dd=1.268, $\Delta$=0.350, pp=0.163, Cohen’s dd=0.736, $\Delta$=0.195, pp=0.571, Cohen’s dd=0.451 for human-75%, 50% and 25% conditions compared with the original AI condition), showing that even limited human involvement improves performance. However, performance decreased steadily as the proportion of AI increased: Human-75% condition (iterations 13-15: M=2.560, $\text{CI}_{95\%}$=[2.242, 2.994]; human-50% (M=2.176, $\text{CI}_{95\%}$=[1.921, 2.467], and human-25% conditions (M=2.021, $\text{CI}_{95\%}$=[1.780, 2.246]), showing a trade-off between the human proportion and the coding performance (rr=-0.986, pp=0.002; Fig. 5C). Moreover, none of the hybrid conditions showed significant improvement with iterations (correlations for human-75%, 50% and 25% were rr=0.043,-0.176, -0.23).
Role division in vibe coding
@@
Beyond the question of whether vibe coding benefit from human-AI hybrid teams, it is an open question how roles should be allocated between different agents across iterations. In our framework, we distinguish between two roles – selection and instruction – reflecting common practices in real-world collaborative programming, where humans are involved both in choosing among alternatives and in providing guidance. To assess the respective contributions of these roles, we conducted an ablation study in which the selection step was removed and the edited output was always propagated to the next iteration.
@@
We reran the experiment using 10 target images, repeated twice, yielding 20 coding chains and a total of 300 iterations. We recruited 30 new human participants to perform the task. As shown in Fig. 6A (light red), removing the selection step produced a significant reduction in human vibe-coding quality relative to the original human condition with selection in the final iteration (iterations 13-15: M=2.331, $\text{CI}_{95\%}$=[2.136, 2.552], which is significant compared with original human condition ($\Delta$=-0.448, pp=0.012, Cohen’s dd=-1.190). This shows that when humans guide the process, explicit evaluation and choice between alternatives helps maintain a coherent direction and supports cumulative improvement, though this effect is relatively small. However, for AI Fig. 6B (cyan), removing the selection had almost no effect (iterations 13-15: M=1.884, $\text{CI}_{95\%}$=[1.586, 2.233], compared with the original AI condition $\Delta$=0.058, pp=1.000, Cohen’s dd=0.112) suggesting that AI-generated guidance already follows a fixed internal preference, and explicitly comparing alternatives does not improve its trajectory.
@@
Figure 6: Role division for collaborative vibe coding. (A) Validation results of experiments removing the code selection part. (B) Validation results of experiments setting human/AI agents in different roles (instructor/selector).
@@
Fig 6B (light red) presents the results of a complementary ablation study with 69 participants in which the AI system assumed the selection role while humans provided the instruction. Replacing the human selector with an AI selector, while retaining a human instructor, yielded performance comparable to fully human-dominated vibe coding (iterations 13-15: M=2.748, $\text{CI}_{95\%}$=[2.552, 2.980], compared with the original human condition $\Delta$=-0.031, pp=1.000, Cohen’s dd=-0.080) suggesting that selection can be offloaded to AI with little loss in performance, offering a practical way to reduce human effort without sacrificing the quality of collaborative vibe coding.
@@
This further suggests that human instructors tend to generate effective improvements. Moreover, human selection becomes essential when the instructor is an AI system (M=2.485, $\text{CI}_{95\%}$=[2.055, 2.930], compared with the original AI condition $\Delta$=0.659, pp=0.025, Cohen’s dd=1.055). Taken together, these findings highlight the importance of human supervision in mitigating compromised or suboptimal AI-generated outputs. Further analysis on selection acceptance rate (whether selector chooses the current SVG) proves this by revealing the misalignment between human and AI when collaborating on a task. Specifically, the acceptance rate when humans act as both instructor and selector (M=0.661, $\text{CI}_{95\%}=[0.620,0.700]$) is significantly higher than the condition with a human instructor and an AI selector (M=0.379, $\text{CI}_{95\%}=[0.334,0.423]$, difference $\Delta$=0.282, pp<0.001, Cohen’s dd=0.589) and vice versa (M=0.342, $\text{CI}_{95\%}=[0.295,0.388]$, $\Delta$=0.320, pp<0.001, Cohen’s dd=0.674).
Robustness to AI model type and social information
@@
To test the robustness of our experimental vibe coding pipeline to other factors, we conducted several control experiments. First, we changed how the AI vibe coder evaluated the SVG outputs. In the main experiment, the AI vibe coder observed the SVG code directly. In the control conditions, we instead allowed the AI vibe coder to view the rendered SVG image only, matching the information available to human vibe coders, or to view both the SVG code and its rendered image. Across both viewing modes, AI-led vibe coding showed similar performance to the main AI condition, remaining significantly worse than human-led vibe coding ($\Delta$=-1.138, -0.545, both pp<0.01) and exhibiting no improvement across iterations (View SVG-rendered image: rr=-0.026, $\text{CI}_{95\%}$=[-0.174, 0.127]; View both SVG code and rendered image: rr=-0.321, $\text{CI}_{95\%}$=[-0.451, -0.181]; Fig. 7A). This indicates that the AI’s failure is not due to limited access to visual information.
@@
We also tested the advanced method from previous research, named FeedBack Descent [20], where the selector not only provided the selection but also gave feedback on the code. Similarly, there is no improvement across iterations (rr=-0.07, $\text{CI}_{95\%}$=[-0.268, 0.110]) and the final performance is significantly lower than the human-led condition ($\Delta$=-0.677, pp=0.019, Cohen’s dd=-1.081; Fig. 7A).
We further tested whether the observed effects depend on the specific AI model used for vibe coding. In addition to GPT-5, we repeated the AI-led experiments using Claude-4.5-Opus (Anthropic, [3]) and Gemini-3-Pro (Google, [38]) as instructor and selector, while keeping the code generator to be GPT-5 (to facilitate a fair comparison between the AI model performance). Both alternative models produced lower final similarity scores than human-led vibe coding (Claude: M=2.323, $\text{CI}_{95\%}$=[1.979, 2.847]; Gemini: M=1.933, $\text{CI}_{95\%}$=[1.617, 2.324]) and showed the similar pattern of no improvement across rounds (Claude: rr=-0.046, $\text{CI}_{95\%}$=[-0.211, 0.117]; Gemini: rr=-0.129, $\text{CI}_{95\%}$=[-0.278, 0.026]; Fig. 7B). These results demonstrate that the performance gap and collapse effect are not specific to a single model, but instead reflect a general limitation of current AI systems when acting as high-level guidance providers in multi-round collaborative coding.
@@
Figure 7: Control experiments for vibe coding pipeline. (A) Experiment results of different viewing modes in AI-led vibe coding. (B) Results of different AI-led vibe coding.
Discussion
@@
Here, we systematically explore human-AI compositions and roles in vibe coding. Across multiple experiments, human-led coding consistently improved over iterations, while AI-led coding often collapsed despite access to the same information and similar execution capabilities. This points to key struggles of today’s AI systems in sustaining coherent high-level direction across repeated interactions, of the kind necessary for successful vibe coding. Our results highlight that studying human–AI collaboration directly reveals behaviors that would be missed by evaluating autonomous AI agents alone.
@@
A key finding is the systematic misalignment between how humans and AI use language to guide collaboration. Humans tend to issue short, action-oriented instructions that focus on what to change next while AI produces long, descriptive instructions that emphasize details. Importantly, constraining instruction length did not eliminate this gap, indicating that misalignment arises from differences in strategy rather than verbosity. This can also be supported by our results that AI prefers its own creations rather than humans’. This aligns with broader concerns that language models may optimize for descriptive completeness instead of goal-directed coordination in interactive tasks [39, 28, 25].
@@
Our hybrid and role-division experiments further clarify how labor should be allocated in human–AI systems. Partial human involvement improved performance relative to fully AI-driven pipelines, but even limited AI guidance could shift the coding trajectory in ways that later human input could not fully correct. At the same time, decomposing the vibe coder role showed that high-level idea generation and instruction are the critical human contributions, whereas evaluation and selection can often be delegated to AI without loss in performance. This suggests a practical design principle for hybrid systems: humans should set direction, while AI can support evaluation and execution.
@@
More broadly, our findings contribute to a cognitive science of hybrid societies by showing that collective performance depends not only on agent competence, but on how roles, feedback, and control are assigned and structured over time [5]. Vibe coding offers a tractable model system for studying these dynamics, with implications beyond programming, toward other creative design work and decision-making in human–AI teams [2, 4, 32, 9, 7].
Limitations
@@
Several limitations of our study should be noted. First, in our task, participants provided improvement instructions only after selecting the better SVG, whereas in real-world coding, people often also comment on the weaknesses of the rejected option; future work should incorporate comparative feedback on unselected outputs. Second, although SVG coding is intuitive and accessible, it remains a simplified task, and future studies should test more realistic coding settings, such as website or game development. Third, we used a single AI model as either a coding assistant, instructor, or selector, while prior work has explored mixtures of different AI systems; whether combining multiple AI agents improves vibe coding remains an open question. Finally, our experiments model collaboration as a linear iterative chain, whereas real-world coding often involves team-based, parallel collaboration, which should be explored in future work. We also note that we only study short-term human-AI interaction effects, but future work should also consider long-term feedback loops and risks like cognitive decline and overreliance [29, 16, 6].
Conclusion
@@
To conclude, in this work, we introduced a controlled experimental framework for studying collaborative vibe coding and used it to systematically compare human- and AI-led vibe coding. We show that humans provide uniquely effective high-level guidance across iterations, that AI guidance often leads to performance collapse. Also, we find that a careful role allocation which keeps humans in charge of direction while offloading evaluation to AI can improve hybrid performance. Together, these results demonstrate the value of studying human–AI collaboration directly and offer concrete insights for designing effective hybrid human–AI systems.
Acknowledgments

### v2

@@ 2.3 Testing hybrid human-AI instructors and selectors
Building on the previous results showing a strong advantage of humans as selectors and instructors, we next asked whether this advantage could be partially retained while reducing human involvement. Human labor is costly and difficult to scale, whereas AI can provide fast and generally less expensive automatic guidance. We next tested a hybrid human–AI vibe coding setting that mixes human and AI vibe coders within the same multi-round coding chain with the same code generation system.
Figure 5: Hybrid Human-AI-Led vibe coding. (A) The schematics of human-AI hybrid vibe coding. (B) Validation rating results of human-AI hybrid vibe coding. (C) Trade-off between human proportion and the final vibe coding performance. Error bars correspond to 95% CI, same as the shaded area in (B).
@@
In this experiment, we constructed hybrid coding chains in which selector and instructor roles were randomly assigned to either a human or an AI agent according to a fixed ratio (Fig. 5A). As in all other experiments, human and AI instructors each interacted with an AI code generation system to create the SVG code. We tested three hybrid conditions: human-75% (AI-25%, where humans take 75% of iterations and AI takes the rest), human-50% (AI-50%), and human-25% (AI-75%).
@@
As shown in Fig. 5B, all hybrid conditions performed better than fully AI-led vibe coding ($\Delta=.55$, $p<.001$, $d=.79$, for human-75%; $\Delta=.34$, $p=.003$, $d=.481$, for human-50%; $\Delta=.17$, $p=.056$, $d=.27$, for human-25% compared with the original AI condition), showing that even limited human involvement improves performance. However, performance decreased steadily as the proportion of AI increased: human-75% condition (iterations 13-15: $M=2.23$, $\text{CI}_{95\%}$=[2.05, 2.43]; human-50%: $M=2.02$, $\text{CI}_{95\%}$=[1.83, 2.22], and human-25% conditions ($M=1.85$, $\text{CI}_{95\%}$=[1.70, 2.01]), showing a trade-off between the human proportion and the coding performance (r=.97r=-.97, $p=.006$; Fig. 5C). Moreover, none of the hybrid conditions showed positive improvement with iterations (correlations for human-75%, 50% and 25% were r=.07r=-.07, .12-.12, .23-.23, respectively).
2.4 Role division in vibe coding
@@
Beyond the question of whether vibe coding generally benefits from human-AI hybrid teams, we next asked how roles should be allocated between different agents across iterations. In our framework, we distinguish between two roles – selectors and instructors – reflecting common practices in real-world collaborative programming, where humans are involved both in choosing among alternatives and in providing guidance. To assess the respective contributions of these roles, we conducted an ablation study in which the selection step was removed and the edited output was always propagated to the next iteration.
@@
We recruited 31 new human participants for 20 coding chains (experiment 10 in Appendix Table 2). As shown in Fig. 6A (light red), removing the selection step produced a significant reduction in human vibe-coding quality relative to the original human condition with selection in the final iteration (iterations 13-15: $M=2.26$, $\text{CI}_{95\%}$=[2.09, 2.42], which is significant compared with original human condition ($\Delta=-.44$, $p<.001$, d=.62d=-.62). This shows that when humans guide the process, explicit evaluation and choice between alternatives help maintain a coherent direction and support cumulative improvement. However, for AI as shown in Fig. 6A (cyan), removing the selection had almost no effect (iterations 13-15: $M=1.68$, $\text{CI}_{95\%}$=[1.56, 1.81], compared with the original AI condition $\Delta=.05$, $p=.339$, $d=.08$) suggesting that AI-generated guidance already follows a fixed internal preference, and explicitly comparing alternatives does not improve its trajectory.
@@
Figure 6: Role division for collaborative vibe coding. (A) Validation results of experiments removing the code selection step. (B) Validation results of experiments setting human/AI agents in different roles (instructor/selector).
@@
Fig. 6B (light red) presents the results of a complementary ablation study with 78 participants in which the AI system assumed the selection role while humans provided the instruction (experiment 12 in Appendix Table 2). Replacing the human selector with an AI selector, while retaining a human instructor, yielded performance comparable to fully human-dominated vibe coding (iterations 13-15: $M=2.88$, $\text{CI}_{95\%}$=[2.68, 3.08], compared with the original human condition $\Delta=.18$, $p=.081$, $d=.23$) suggesting that selection can be offloaded to AI with little loss in performance, offering a practical way to reduce human effort without sacrificing the quality of collaborative vibe coding.
@@
This further suggests that human instructors tend to generate effective improvements. Moreover, human selection becomes essential when the instructor is an AI system (experiment 13 in Appendix Table 2, $M=2.27$, $\text{CI}_{95\%}$=[1.96, 2.59], compared with the original AI condition $\Delta=.59$, $p<.001$, $d=.77$). Taken together, these findings highlight the importance of human supervision in mitigating compromised or suboptimal AI-generated outputs. Further analysis of the selection acceptance rate (i.e., whether the selector chooses the current SVG) supports this by revealing misalignment between humans and AI when collaborating on a task. Specifically, the acceptance rate when humans act as both instructor and selector ($M=.66$, $\text{CI}_{95\%}=[.62,.70]$) is significantly higher than the condition with a human instructor and an AI selector ($M=.38$, $\text{CI}_{95\%}=[.33,.42]$, difference $\Delta=.28$, $p<.001$, $d=.59$) and vice versa ($M=.34$, $\text{CI}_{95\%}=[.30,.39]$, $\Delta=.32$, $p<.001$, $d=.67$).
2.5 Robustness to AI model type and social information
@@
To test the robustness of our experimental vibe coding pipeline to other factors, we conducted several control experiments. First, we changed how the AI vibe coder evaluated the SVG outputs (Fig. 7A). In the main experiment, the AI vibe coder observed the SVG code directly. In the control conditions, we instead allowed the AI vibe coder to view the rendered SVG image only, matching the information available to human vibe coders, or to view both the SVG code and its rendered image. Across both viewing modes, AI-led vibe coding showed similar performance to the main AI condition, remaining significantly worse than human-led vibe coding ($\Delta=-.59$, 1.25-1.25, both $p<.001$) and exhibiting no improvement across iterations (Only SVG-rendered image visible: r=.01r=-.01, $\text{CI}_{95\%}$=[-.17, .15]; Both SVG code and rendered image visible: r=.32r=-.32, $\text{CI}_{95\%}$=[-.46, -.18]). This indicates that the AI’s failure is not due to limited access to visual and code information.
@@
We also tested an advanced recent AI pipeline, named FeedBack Descent [lee2025feedback], where the selector not only provided the selection but also gave feedback on the code (for more details, see Appendix A.2.4). Similarly, there was no improvement across iterations (r=.08r=-.08, $\text{CI}_{95\%}$=[-.26, .09]) and the final performance is significantly lower than the human-led condition ($\Delta=-.88$, $p<.001$, d=1.32d=-1.32; Fig. 7A).
@@
Figure 7: Control experiments for vibe coding pipeline. (A) Experiment results of different viewing modes in AI-led vibe coding, with GPT-5 acting as instructors and selectors. (B) Results of AI-led vibe coding, under different AI systems.
We further tested whether the observed effects depend on the specific AI model used for vibe coding (Fig. 7B). In addition to GPT-5, we repeated the AI-led experiments using Claude-4.5-Opus (Anthropic, [anthropic2025opus]) and Gemini-3-Pro (Google, [team2023gemini]) as instructor and selector, while keeping the code generator to be GPT-5 (to facilitate a fair comparison between the AI model performance). Both alternative models produced lower final similarity scores than human-led vibe coding (Claude: $M=2.25$, $\text{CI}_{95\%}$=[1.93, 2.59]; Gemini: $M=1.81$, $\text{CI}_{95\%}$=[1.63, 2.01]) and showed a similar pattern of no improvement across rounds (Claude: $r=.02$, $\text{CI}_{95\%}$=[-.14, .18]; Gemini: r=.08r=-.08, $\text{CI}_{95\%}$=[-.23, .07]; Fig. 7B). These results demonstrate that the performance gap and collapse effect are not specific to a single model, but instead reflect a general limitation of current AI systems when acting as high-level guidance providers in multi-round collaborative coding. As the performance gap was getting somewhat smaller with a more advanced AI model (Claude vs. human-led: $\Delta=-.45$, $p=.005$, d=.56d=-.56), our results also indicate that the improvement of AI may enhance the alignment between AI and human behavior.
3 Discussion
@@
As vibe coding becomes an increasingly widespread practice, it is important to characterize how it impacts collective outcomes, and what arrangements of human-AI collaboration yield the best outcome. Our experimental paradigm makes it possible to systematically explore different human-AI compositions and roles in vibe coding. Across 18 experiments (and two evaluation experiments), human-led coding consistently improved over iterations, while AI-led coding often collapsed despite access to the same information and similar execution capabilities. This highlights significant challenges for the AI systems we studied in maintaining coherent high-level cumulative improvements across repeated interactions, a critical factor for successful vibe coding. Our results show how studying human–AI collaboration directly reveals behaviors that would be missed by evaluating autonomous AI agents alone.
@@
One of our key findings is the systematic misalignment between how humans and AI systems use language to guide collaboration. Humans tend to provide short, action-oriented instructions that focus on what requires changing, while AI systems produce long, descriptive instructions that emphasize details. Importantly, constraining instruction length did not eliminate this gap, indicating that misalignment arises from differences in strategy rather than verbosity. This is also supported by our results that AI systems have lower sensitivity to the quality difference between human- and AI-led vibe coding creations, alongside a tendency to evaluate their own creations more favorably. This aligns with broader concerns that large language models may optimize for descriptive completeness instead of goal-directed coordination in interactive tasks [tsvilodub2023overinformative, miehling2024language, ma2025pragmatics].
@@
Our hybrid and role-division experiments further clarify how labor could be effectively allocated in human–AI systems. Partial human involvement improved performance relative to fully AI-driven pipelines, but even limited AI guidance could shift the coding trajectory in ways that later human input could not fully correct. At the same time, decomposing the vibe coder role showed that high-level idea generation and instruction are – in our setup – the critical human contributions, whereas evaluation and selection can often be delegated to AI without loss in performance. This suggests a practical design principle for hybrid systems: humans should set direction, while AI can support evaluation and execution.
Advances in contemporary AI systems are proceeding at an extraordinary pace. In principle, coding agents specifically optimized for tasks such as SVG image generation can achieve high levels of performance. Indeed, in our experiments, the model we evaluated matched human participants at the outset (iteration 1). Much of current engineering efforts are directed toward maximizing performance at this initial stage. However, when performance is examined at the level of group dynamics across successive iterations, a different pattern emerges: AI-led systems tend to deteriorate over time rather than improve. Although further task-specific optimization – such as models tailored explicitly for SVG generation – may substantially reduce or even eliminate the overall performance gap relative to humans, important insight of our work concerns the critical contrast between strong initial performance and subsequent degradation. As vibe coding becomes more widespread and is deployed for increasingly complex tasks, understanding and addressing this divergence will be essential, highlighting the broader implications of our findings.
@@
More broadly, our findings contribute to a cognitive science of hybrid societies by showing that collective performance depends not only on agent competence, but on how roles, feedback, and control are assigned and structured over time between humans and machines [bhatt2025should]. Vibe coding offers a tractable model system for studying these dynamics, with implications beyond programming, toward other creative design work and decision-making in human–AI teams [amershi2019guidelines, bansal2021does, rezwana2023designing, davis2025co, collins2024building].
3.1 Limitations
@@
Several limitations of our study should be noted. First, in our task, participants provided improvement instructions only after selecting the preferred SVG, whereas in real-world coding contexts, developers typically also articulate the shortcomings of the rejected alternative. Initial tests implementing a recent feedback-based method from the literature [lee2025feedback] did not yield substantial performance gains (Figure 7). Nevertheless, future work should more systematically explore alternative feedback paradigms; for example, structured code-review procedures modeled on open-source development practices may provide richer supervisory signals. Second, although SVG coding is intuitive and accessible, it remains a simplified task, and future studies should test more realistic coding settings, such as website or game development. Third, we used a single AI model as either a coding assistant, instructor, or selector, while our prior work has explored mixtures of different AI systems [shiiku2025dynamics]; whether combining multiple AI agents improves vibe coding remains an open question. Finally, our experiments model collaboration as a linear iterative chain, whereas real-world coding often involves team-based, parallel collaboration, which should be explored in future work [shiiku2025dynamics, marjieh2025characterizing]. We also note that we only study short-term human-AI interaction effects, but future work should also consider long-term feedback loops and risks like cognitive decline and overreliance [oktar2025identifying, ibrahim2025measuring, collins2025revisiting].
3.2 Conclusion
@@
Understanding the potential impacts of AI systems on human society is going to require empirical work that explores the behavior of human-AI hybrid systems. In this work, we introduced a controlled experimental framework for studying collaborative vibe coding and used it to systematically compare human- and AI-led processes. We show that humans provide uniquely effective high-level guidance across iterations and that AI guidance often leads to performance collapse. Additionally, we find that a role allocation that keeps humans in charge of direction while offloading evaluation to AI can improve hybrid performance. Together, these results demonstrate the value of studying human–AI collaboration directly and offer concrete insights for designing effective hybrid human–AI systems.
4 Acknowledgments